THE PURE ROTATIONAL SPECTRUM OF THE CoS RADICAL (X^{4}Δ_{i}): CONTINUED STUDIES IN MONO-LIGATED COBALT SPECIES

\underline{M. A. FLORY}, S. K. McLAMARRAH, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Obsevatory, University of Arizona, Tucson, AZ 85721.

The sub-millimeter wave spectrum of the CoS radical has been observed in the frequency range 350-540 GHz using direct absorption techniques. This work is the first spectroscopic study of this molecule. Cobalt vapor was reacted with CS_{2} to produce CoS. Rotational transitions from three spin-components in the radical’s X ^{4}Δ_{i} ground state have been recorded thus far, and magnetic hyperfine splittings due to the ^{59}Co nucleus (I = 7/2) have been resolved. We are continuing our search to identify the fourth spin-component. Transitions from the Co^{34}S isotopomer were also measured, as well as several arising from excited vibrational levels. There is no evidence of λ -doubling in the lower three spin-components (Ω = 7/2, 5/2, 3/2). The hyperfine pattern in the various spin-orbit components is somewhat irregular, particularly the Ω = 3/2, suggesting some mixing from nearby excited electronic states. The preliminary data have been fit with a case (a) Hamiltonian, and spectroscopic constants have been determined. The most current results of this study will be presented.